Smullyan Rudimentary Predicates and Skolem Elementary Functions on Trees

Anatoly P. Beltiukov

Analogs of Smullyan rudimentary predicates and Skolem elementary functions on binary trees are constructed. Tree analogs of Grzegorczyk classes based on bounded recursion are also considered. Differences between the properties of these classes and those of ordinary Smullyan, Skolem and Grzegorczyk classes are studied.

Address: Mathematics Department, Udmurt State University, Izhevsk, 426034, Russia

E-mail: belt@uni.udm.ru